In a solid, the speed of elastic longitudinal wave is 1.35ms -1. If the Young’s modulus  of
elasticity of the solid is 2X1011 N m-2  × calculate its mass density.

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Speed of longitudinal wave is given asv=\sqrt{\dfrac{Y}{\rho}}
\implies \rho=\dfrac{Y}{v^2}=\dfrac{2\times 10^{11}}{1.35^2}kg/m^3=11\times 10^{10}kg/m^3